OpenStudy (anonymous):

(cos Θ + cos Θ)^2 + (cos Θ + cos Θ)^2

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Why is this troubling? Are they ALL \(\cos(\theta)\)? Add them inside the parentheses and see what happens.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

this is like \[(a+a)^2+(a+a)^2\]

OpenStudy (anonymous):

But the answers are: sin^2 Θ 4 8cos^2 Θ cos^2 Θ

OpenStudy (anonymous):

you get \[(2a)^2+(2a)^2=4a^2+4a^2=8a^2\]

OpenStudy (anonymous):

Ohhh so C

OpenStudy (anonymous):

ok there is no trig in this it is algebra only

OpenStudy (anonymous):

What do you mean by that @satellite73